# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 4.3. Preliminary results
where we use (4.7) in the last step. The upper bound is finite by the conditions in (PHC). We thus arrive at (4.92). This implies that the first line of the right-hand side of (4.90) is $\mathbb{P}_{\mathcal{S}}$-almost surely finite conditionally on $\mathcal{F}_{T}$. We proceed with the sets on the second line. We show that for any $u\in\mathcal{A}(K)\setminus\{V_{1}\}$ such that $D^{(u)}=\infty$,
[…]
@@
[…]
The first probability being equal to one $\mathbb{P}_{\mathcal{S}}$-almost surely follows directly from Proposition 4.13. Indeed, by the definition of $V_{1}$ and $V_{2}$, it follows that any individual $u\in\mathcal{A}(K)\setminus\{V_{1}\}$ has given birth to fewer than $K$ children by time $T$ if $u\neq V_{2}$ and to exactly $K$ children if $u=V_{2}$. Hence, the event in the conditional probability is independent of $\mathcal{F}_{T}$. Proposition 4.13 then yields the desired result, by replacing $\varnothing$ with $u$ and ‘$v$ catches up to $\varnothing$’ with ‘$uv$ catches up to $u$’.
The proof of (4.104) follows in an analogous way when following the proof of Proposition 4.13, since we know that $V_{2}$ has produced $K$ children by time $T$, whereas the individuals $uv_{1}$ with $v_{1}>K$ have all not been born by time $T$.
To conclude the proof, it remains to show that the set on the final line on of (4.90) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$. That is, for $j\in\{1,2\}$,
[…]
Here, individuals $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants may have been born by time $T$, so that a slightly more careful argument is required. Let us define
$\mathcal{T}_{T}:=\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)\leq T\}$ (4.106)
to be the sub-tree rooted at $V_{1}$ but restricted to the children $V_{1}(K+1),V_{1}(K+2),\ldots$ and their descendants that are born by time $T$. It is clear that $\mathcal{T}_{T}\subset\mathrm{BP}(T)$, so that it is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, and that it is $\mathcal{F}_{T}$ measurable. For $w\in\mathcal{T}_{T}$, we define
$K(w):=\max\{0\leq k\leq D^{(w)}:wk\in\mathcal{T}_{T}\}$ (4.107)
to be the ‘last’ child of $w$ that is included in $\mathcal{T}_{T}$. We then have the inclusion
$\{V_{1}v:{}$ $v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{j}\}$ (4.108)
$\subseteq\bigcup_{w\in\mathcal{T}_{T}}\{wv:v\in\mathcal{U}_{\infty},v_{1}>K(w),D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{j}\}.$
As a result, we have
$\mathbb{E}_{\mathcal{S}}\Big{[}{}$ |$|\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.109)
[…]
Each expected value in the sum on the right-hand side can be written as
$\sum_{r=1}^{\infty}\sum_{v_{1}=K(w)+1}^{D^{(w)}}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(wv)}=\infty,wv\text{ c.u.t.\ }V_{j}\,\Big{|}\,\mathcal{F}_{T}\right),$ (4.110)
and each such probability equals
$\mathbb{P}_{\mathcal{S}}\bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\mathcal{B}(wK(w))+\!\!\!\!\sum_{i=K(w)+1}^{v_{1}}\!\!\!\!X(wi)+\sum_{\ell=2}^{r}\sum_{i=1}^{v_{\ell}}X(wv_{1}\cdots v_{\ell-1}i)+\!\!\!\!\sum_{i=1}^{K_{1}+n}\!\!X(wvi)$ (4.111)
$\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!X(V_{j}i),D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,r\}\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
We observe that
$\mathcal{B}(wK(w))$ $<T<\mathcal{B}(w(K(w)+1))=\mathcal{B}(wK(w))+X(w(K(w)+1)),$
and
$\mathcal{B}(V_{j}K_{j})$ $\leq T<\mathcal{B}(V_{j}(K_{j}+1))=\mathcal{B}(V_{j}K_{j})+X(V_{j}(K_{j}+1)),$
with equality in the second line only for $j=2$. Moreover, again using the memoryless property of the exponential distribution,
$\mathcal{B}(wK(w))+X(w(K(w)+1))-T$ $\sim X(w(K(w)+1))^{\prime}$
and
$\mathcal{B}(V_{j}K_{j})+X(V_{j}(K_{j}+1))-T$ $\sim X(V_{j}(K_{j}+1))^{\prime},$
where $X(w(K(w)+1))^{\prime}$ and $X(V_{j}(K_{j}+1))^{\prime}$ are i.i.d. copies of $X(w(K(w)+1))$ and $X(V_{j}(K_{j}+1))$, respectively, independent of all other random variables and also independent of $\mathcal{F}_{T}$. We thus obtain that the conditional probability can be bounded from above by
$\mathbb{P}_{\mathcal{S}}\bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\sum_{i=K(w)+1}^{v_{1}}\!\!\!\!X(wi)+\sum_{\ell=2}^{r}\sum_{i=1}^{v_{\ell}}X(wv_{1}\cdots v_{\ell-1}i)+\!\!\!\!\sum_{i=1}^{K_{j}+n}\!\!X(wvi)\leq\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!X(V_{j}i),$ (4.112)
$D^{(wv)}=\infty,D^{(wv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in\{2,\ldots,r\}\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
The dependence on $\mathcal{F}_{T}$ is only in the random variables $K(w)$ and $K_{j}$, all other random variables are independent of $\mathcal{F}_{T}$. Also using that the offspring random variables are independent of the inter-birth random variables and applying Lemma 4.8 thus yields the upper bound
$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K_{j}+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\prod_{i=K(w)+1}^{v_{1}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.113)
×$\times\prod_{\ell=2}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$
Using this in (4.110) yields
$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K_{j}+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.114)
×$\times\sum_{r=1}^{\infty}{}$ $\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\prod_{\ell=2}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$
Since $K_{j}\geq K$ by definition, we obtain the upper bound
$\mathbb{P}\!\left(D=\infty\right){}$ $\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\sum_{v_{1}=K(w)+1}^{D^{(w)}}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.115)
×$\times\sum_{r=0}^{\infty}\Bigg{(}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v}X(i)\bigg{)}\right]\Bigg{)}^{r}$
[…]
where we use (4.7) in the final step. By the conditions in (PHC), the first three terms are finite. The final sum is finite when $D^{(w)}$ is finite. In the case that $D^{(w)}=\infty$, it equals
$\sum_{v_{1}=K(w)+1}^{\infty}\prod_{i=K(w)+1}^{v_{1}}\!\!\!\!\!\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ $\leq\mathbb{P}\!\left(D=\infty\right)^{-1}\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]^{-1}\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\prod_{i=1}^{v}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]$ (4.116)
$=\mathbb{P}\!\left(D=\infty\right)^{-1}\prod_{i=1}^{K(w)}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]^{-1}\widehat{\mu}(\lambda),$
where we again use (4.7) to obtain the final equality. This upper bound is $\mathbb{P}_{\mathcal{S}}$-almost surely finite for any $w\in\mathcal{T}_{T}$, since $K(w)<\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, since $\mathcal{T}_{T}$ is a set wit a $\mathbb{P}_{\mathcal{S}}$-almost surely finite size, the upper bound in (4.109) is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, which yields (4.105).
Combined with (4.104), (4.103), and (4.92), we thus conclude that
$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{P}_{2}|<\infty\,|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\mathrm{-a.s.}$ (4.117)
Since the probability measure is bounded from above, it follows that $\mathcal{P}_{2}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, which completes the proof. ∎
Adaptations for the proof of Proposition 4.4 with $m>2$. We now discuss the adaptations of the proof for $m=2$ to extend to the general case $m\geq 2$. To show that $\mathcal{P}_{m}$ has a finite size when $m>2$, we define
$T:=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}\geq m\},$ (4.118)
and we let $V_{1},\ldots,V_{m}$ be the first $m$ individuals that produce $K$ children, ordered such that $\mathcal{B}(V_{1}K)<\mathcal{B}(V_{2}K)<\cdots<\mathcal{B}(V_{m}K)=T$. The sets $\mathcal{A}(K)$ and $\mathcal{A}^{+}$ have the same definition. We then have
$\mathcal{P}_{m}\subseteq\mathrm{BP}(T){}$ $\cup\bigcup_{u\in\mathcal{A}^{+}}\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1},V_{2},\ldots,\text{ or }V_{m}\}$ (4.119)
$\cup\bigcup_{u\in\mathcal{A}(K)\setminus\{V_{1},\ldots,V_{m-1}\}}\!\!\!\!\!\!\!\!\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{c.u.t.\ }u,\text{ or }V_{2},\ldots,V_{m}\}$
$\cup\bigcup_{j=1}^{m-1}\{V_{j}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{j}v)>T,D^{(V_{j}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{1},\ldots,\text{ or }V_{m}\}.$
4.5. Unique persistent $m$-hubs
We conclude this section by proving Theorem 4.5.
Proof of Theorem 4.5.
It follows from Theorem 4.2 that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when Assumption 4.1 and (PHC) are satisfied. From the proof of Theorem 4.2 (using Proposition 4.14 in particular), it follows that the number of persistent $m$-hubs is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. We assume that there is more than one persistent $m$-hub with positive probability and argue by contradiction. Suppose that both $u$ and $v$ are both persistent $m$-hubs. It follows that there exists $T>0$ such that their offspring is the same for all time $t\geq T$. That is, $\deg_{t}(u)=\deg_{t}(v)$ for all $t\geq T$. This is equivalent to there existing $N\in\mathbb{N}$ such that $\mathcal{B}(uk)=\mathcal{B}(vk)$ for all $k\geq N$. That is,
$\mathcal{B}(u)+\sum_{i=1}^{k}X(ui)=\mathcal{B}(v)+\sum_{i=1}^{k}X(vi)\qquad\text{or, equivalently,}\qquad\mathcal{B}(u)-\mathcal{B}(v)+\sum_{i=1}^{k}(X(ui)-X(vi))=0.$ (4.120)
Let us set $S_{i}:=X(ui)-X(vi)$. It is clear that the series $\sum_{i=1}^{\infty}S_{i}$ is finite almost surely by Assumption 4.1 and the second condition in (PHC). Since $\mathcal{B}(v)$ and $\mathcal{B}(u)$ can be represented as finite sums of inter-birth random variables, the equality on the right can only hold with positive probability when the distribution of $\sum_{i=1}^{\infty}S_{i}$ has an atom. However, the conditions in Theorem 4.5, combined with Lemma 4.9 and Remark 4.10, yield that the distribution does not have an atom. This contradiction thus implies that $u=v$ $\mathbb{P}_{\mathcal{S}}$-almost surely and there thus exists a unique persistent $m$-hub.
To conclude the proof, we show (4.5). We know that there exists a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$. Suppose $u$ is the unique persistent $m$-hub and $v$ is the unique persistent $m^{\prime}$-hub, with $1\leq m<m^{\prime}<\infty$ two integers. We now assume that
$\mathbb{P}_{\mathcal{S}}\left(\exists T>0\ \forall t\geq T:\deg_{t}(u)>\deg_{t}(v)\right)=0,$ (4.121)
and aim to show a contradiction. Since $u$ and $v$ are the unique $m$-hub and $m^{\prime}$-hub, respectively, with $m<m^{\prime}$, we have
$\mathbb{P}_{\mathcal{S}}\left(\exists T^{\prime}>0\ \forall t\geq T^{\prime}:\deg_{t}(u)\geq\deg_{t}(v)\right)$ $=1,$
and
$\mathbb{P}_{\mathcal{S}}\left(\exists T^{\prime\prime}>0\ \forall t\geq T^{\prime\prime}:\deg_{t}(u)=\deg_{t}(v)\right)$ $=0.$
Combining this with our assumption, it follows that $\mathbb{P}_{\mathcal{S}}$-almost surely,
$\mathcal{B}(uk)\leq\mathcal{B}(vk)<\mathcal{B}(u(k+1))\text{ holds for infinitely many }k\in\mathbb{N}$ (4.122)
is satisfied. That is, $u$ and $v$ have the same number of offspring in the non-empty intervals $[\mathcal{B}(vk),\mathcal{B}(u(k+1)))$, for infinitely many $k$. This implies that
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{k}(X(vi)-X(ui))<X(u(k+1))\text{holds for infinitely many }k\in\mathbb{N}.$ (4.123)
We now fix an arbitrary $\varepsilon>0$. Under the condition in (4.4), it follows that $X(u(k+1))\leq\varepsilon$ for all but finitely many $k$. Hence, the inequalities
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{k}(X(vi)-X(ui))<\varepsilon$ (4.124)
hold for infinitely many $k$. Then, since the infinite sum $\sum_{i=1}^{\infty}(X(vi)-X(ui))$ is almost surely convergent under the second condition in (PHC) and $(2)$ in Assumption 4.1,
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{\infty}(X(vi)-X(ui))\leq\varepsilon,$ (4.125)
with positive probability. As $\varepsilon$ is arbitrary, it follows that
$\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=1}^{\infty}(X(vi)-X(ui))=0$ (4.126)
holds with positive probability. However, combining the conditions in (4.3) with Lemma 4.9 yields that the distribution of $\sum_{i=1}^{\infty}(X(vi)-X(ui))$ has no atoms. Since $\mathcal{B}(v)$ and $\mathcal{B}(u)$ can be represented as finite sums of inter-birth random variables, it thus follows that (4.126) occurs with probability zero, which leads to a contradiction and concludes the proof. ∎
5. Persistence in the CTPAVD branching process
In this section we prove Theorem 2.6. Unlike in the analysis carried out in Section 4, a persistent $m$-hub, as in the sense of (PH-$m$), does not emerge $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$, since individuals live an almost surely finite time (i.e. $L<\infty$ a.s.). This requires a more direct approach where we provide precise asymptotic expansions of the random variables $O_{t}^{\mathrm{ct}}$ and $I_{t}^{(1),\mathrm{ct}}$ to show that their difference is a tight sequence of random variables (with respect to $t$), as in (3.14).
@@ 5.1. Preliminary results
We first state some preliminary results that are used throughout the remainder of this section. We recall the definition of the CTPAVD branching process in Section 3, and that we write $D$ for the offspring of an individual, $S_{k}$ for the time an individual needs to give birth to $k$ children, and $L:=S_{D+1}$ for the lifetime of an individual. Furthermore, we recall the sequences $\varphi_{1},\varphi_{2},$ and $\rho_{1}$ from (2.9) and define the sequence
$\rho_{2}(k):=\sum_{i=0}^{k-1}\Big{(}\frac{d(i)}{b(i)+d(i)}\Big{)}^{2}\qquad\text{for }k\in\mathbb{N}.$ (5.1)
@@
We then introduce the following notation. First, recall that $\mathcal{S}$ denotes the event that the branching process survives, and recall that
$\mathbb{P}_{\mathcal{S}}\left(\cdot\right):=\mathbb{P}\!\left(\cdot\,|\,\mathcal{S}\right)\quad\text{and}\quad\mathbb{E}_{\mathcal{S}}[\cdot]:=\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ (5.7)
@@
$\mathcal{B}(s,t):=\{u\in\mathcal{U}_{\infty}:\sigma_{u}\in[s,t]\}$ (5.8)
@@
^$\widehat{\mathcal{R}}^{\lambda^{*}}(t):=\int_{0}^{t}\mathrm{e}^{-\lambda^{*}u}\,\mathcal{R}(\mathrm{d}u),$ (5.10)
@@
Proposition 5.3 is a direct result of [32, Theorem $5.4$] and [18] (see also [32, Proposition $1.1$] for a condensed version). We have the following useful corollary, which is a part of Corollary $6.6$ in [23].
@@
Suppose that $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied, and let $\mathcal{S}:=\{\lim_{t\to\infty}|\mathrm{BP}(t)|=\infty\}$ be the event that the process $\mathrm{BP}$ survives. Then,
@@
    Suppose that Assumption (N-E) is satisfied and that $\overline{d}:=\limsup_{i\to\infty}d(i)<\infty$. Then, $\limsup_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\leq\overline{d}$. In particular, when $d$ converges to zero, then $\rho_{1}=o(\varphi_{1})$. Similarly, if $\underline{d}:=\limsup_{i\to\infty}d(i)>0$, then $\liminf_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\geq\underline{d}$. In particular, if $d$ tends to infinity, then $\varphi_{1}=o(\rho_{1})$.

### v2

@@ 4.2. Preliminary results
We can thus use the inclusion for $\mathcal{P}_{2}$ in (4.67), combined with having shown that all sets on the right-hand side of (4.68), (4.69), and (4.70) are finite $\mathbb{P}_{\mathcal{S}}$-almost surely conditionally on $\mathcal{F}_{T}$, to finally arrive at
$\mathbb{P}_{\mathcal{S}}\left(|\mathcal{P}_{2}|<\infty\,|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (4.111)
By taking expectation with respect to $\mathcal{F}_{T}$, we thus arrive at the desired result and conclude the proof. ∎
Adaptations for the proof of Proposition 4.8 with $m=1$ and $m>2$. When $m=1$, the proof simplifies somewhat. In this case, we instead define the stopping time $T$, as in (4.64), as the time at which there is at least $1$ individual $u$ such that $D^{(u)}=\infty$ that has produced at least $K$ children by time $T$. As the inter-birth times are exponentially distributed, it follows that this individual, say $V_{1}$, has produced exactly $K$ children at time $T$. The domination of $\mathcal{P}_{1}$ is similar as in (4.67), though now we no longer need the second line (as $V_{2}$ does not exist in this case). Furthermore, in the first line we require the individuals $uv$ to catch up to $u$ only, and in the third line to $V_{1}$ only. Bounding the set on the first line then directly follows from Proposition 4.8, and bounding the sets on the third line follows in an analogous way to bounding the sets on the right-hand side of (4.70).
When $m>2$, we define $T$ as the first time that $m$ many individuals $V_{1},\ldots,V_{m}$ such that $D^{(V_{i})}=\infty$ have produced at least $K$ children. We bound $\mathcal{P}_{m}$ in the same way as in (4.67), namely
$\mathcal{P}_{m}\subseteq\mathcal{A}_{-}{}$ $\cup\!\!\!\!\!\!\!\!\!\!\bigcup_{\begin{subarray}{c}u\in\mathcal{A}_{-}\setminus\{V_{2},\ldots,V_{m}\}\\ D^{(u)}=\infty\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\{uv\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }u\text{ or }V_{i}\text{ for some }i\in[m]\setminus\{1\}\}$ (4.112)
$\cup\bigcup_{j=2}^{m}\{V_{j}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{i}\text{ for some }i\in[m]\}$
$\cup\bigcup_{u\in\mathcal{A}_{+}}\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{i}\text{ for some }i\in[m]\}.$
5. Persistence in the CTPAVD branching process
In this section we prove Theorem 2.6. Unlike in the analysis carried out in Section 4, a persistent $m$-hub, as in the sense of (PH-$m$), does not exist $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$, since individuals live an almost surely finite time (i.e. $L<\infty$ a.s.). Indeed, an almost surely finite lifetime implies that $O_{t}^{\mathrm{ct}}$ diverges to infinity with $t$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Since $I_{t}^{(1),\mathrm{ct}}\geq O_{t}^{\mathrm{ct}}$ for all $t$, we thus also have $I_{t}^{(1),\mathrm{ct}}\to\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. To show instead that persistence occurs, in the sense of (P), we use a more quantitative approach where we provide precise asymptotic expansions of the random variables $O_{t}^{\mathrm{ct}}$ and $I_{t}^{(1),\mathrm{ct}}$ to show that their difference is a tight sequence of random variables (with respect to $t$), as in (3.19). That is, for a mapping $s\mapsto f(s)$ to be determined, we show that for any $\varepsilon>0$ small and $A>0$ large, there exist $t^{\prime},K,M>0$ such that for all $t\geq t^{\prime}$, the events
$|O_{s}^{\mathrm{ct}}-f(s)|$ $<K\quad\text{for all }s\in[t-A,t+A],$ (5.1)
@@ 5. Persistence in the CTPAVD branching process
$I_{s}^{\mathrm{ct}}-O_{s}^{\mathrm{ct}}$ $<K\quad\text{for all }s\in[t-A,t+A],$ (5.2)
occur with probability at least $1-\varepsilon$. We first collect several preliminary results in Section 5.1 and prove (5.1) in Section 5.2. We then prove several results on the offspring of ‘young’ and ‘old’ individuals in Section 5.3, to finally prove (5.2) in Section 5.4.
@@ 5.1. Preliminary results
We first state some preliminary results that are used throughout the remainder of this section. We recall the definition of the CTPAVD branching process in Section 3, and that we write $D$ for the offspring of an individual, $S_{k}$ for the time an individual needs to give birth to $k$ children, and $L\coloneq S_{D+1}$ for the lifetime of an individual. Furthermore, we recall the sequences $\varphi_{1},\varphi_{2},$ and $\rho_{1}$ from (2.9) and define the sequence
$\rho_{2}(k)\coloneq\sum_{i=0}^{k-1}\Big(\frac{d(i)}{b(i)+d(i)}\Big)^{2}\qquad\text{for }k\in\mathbb{N}.$ (5.3)
@@
The lemma follows from an asymptotic expansion of the tail distribution of $D$, as in (2.6).
@@
Recall that $\mathcal{S}$ denotes the event that the branching process survives, and that
$\mathbb{P}_{\mathcal{S}}\left(\cdot\right)\coloneq\mathbb{P}\!\left(\cdot\,|\,\mathcal{S}\right)\quad\text{and}\quad\mathbb{E}_{\mathcal{S}}[\cdot]\coloneq\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ (5.9)
@@
$\mathcal{B}(s,t)\coloneq\{u\in\mathcal{U}_{\infty}:\mathcal{B}(u)\in[s,t]\}$ (5.10)
@@
$\widehat{\mathcal{R}}^{\lambda^{*}}(t)\coloneq\int_{0}^{t}\mathrm{e}^{-\lambda^{*}u}\,\mathcal{R}(\mathrm{d}u),$ (5.12)
@@
Proposition 5.3 is a direct result of [30, Theorem $5.4$] and [18] (see also [30, Proposition $1.1$] for a condensed version).
@@
Suppose that $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied, and let $\mathcal{S}\coloneq\{\lim_{t\to\infty}|\mathrm{BP}(t)|=\infty\}$ be the event that the process $\mathrm{BP}$ survives. Then,
@@
    Suppose that Assumption (N-E) is satisfied and that $\overline{d}\coloneq\limsup_{i\to\infty}d(i)<\infty$. Then, $\limsup_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\leq\overline{d}$. In particular, when $d$ converges to zero, then $\rho_{1}=o(\varphi_{1})$. Similarly, if $\underline{d}\coloneq\limsup_{i\to\infty}d(i)>0$, then $\liminf_{k\to\infty}\rho_{1}(k)/\varphi_{1}(k)\geq\underline{d}$. In particular, if $d$ tends to infinity, then $\varphi_{1}=o(\rho_{1})$.